OpenStudy (anonymous):

(2squareroot 7- squareroot3)^2

OpenStudy (anonymous):

(a-b)^2 = a^2 -2ab + b^2 alright so you have \[(2\sqrt{7} - \sqrt{3})^{2}\] = \[28 - 4\sqrt{21} + 3\] = \[-4\sqrt{21} + 31\]